solve
x+4y=19
5x+y+0

solve
a-b=25
3a+4b=4

Answer by texttutoring(324)

x+4y=19
5x+y+0
Let's solve this one by elimination method. Multiply the second equation by 4, then subtract it from the first equation:

x+4y = 19
20x +4y = 0
-----------
-19x = 19

x= -1

Plug this value into either equation to find y:
5x+y = 0
5(-1) + y = 0
-5 + y = 0
y = 5

So the solution is (-1,5) or x=-1, y=5.

Second question is weird because it has no nice answers. Are you supposed to write your answers as fractions or as decimals? I did it with fractions:

a-b=25
3a+4b=4
Multiply the first equation by 3, then subtract.

3a - 3b = 75
3a + 4b = 4
--------------
0 - 7b = 71

b = -71/7

Plug this into either equation to find a:

3a - 3(-71/7) = 75
3a + 213/7 = 75
3a = 75- 213/7
3a = 312/7
a = 312/21
a = 104/7